11. A proton (+q) is 50nm away from an electron (-q) which is 30nm away from another proton (+2q) with

a charge +2q. What is the net electric force on +2q?

12. What is the direction of the net electric force on +2q?

13. What is the net electric field at +2q?

14. What is the net electric force on the proton (+q) in number 11?


1
Expert's answer
2022-03-07T11:44:56-0500

"F=2qk(\frac{q}{r_1^2}-\frac q{r_2^2})=3.5~\mu N,"

"E=2qk(\frac{q}{r_1}-\frac q{r_2})=5.4~\frac{ pN}C."